radomes.openskiesproject.org
BrainiestStudent3458
BrainiestStudent3458 BrainiestStudent3458
  • 06-12-2016
  • Mathematics
Answered

If the probability of a win is 0.24 and the probability of a draw is 0.16, what is the probability of a loss

Verified Answer :

MajorStealth
MajorStealth MajorStealth
  • 06-12-2016

The probability of a loss is .60

.24+.16=.40
1.00-.40=.60

Answer Link

Related Questions

can 21 over 42 be reduced
o Guten Tag Was kann ich 1 fSie tun 2 W würden gern zusammen Urlaub 3 m Ich möchte mich 4 er und ausruhen am 5 lie o Und das wäre an der Ostsee 6 Ada gibt es ei
red roses cost 3 each Pink roses cost 2 each A man buys 24 flowers for his wife with some pink roses and the rest red roses The total cost of the flowers is 68
Why do you need to place a second tube into the centrifuge
solve the system of equations below by graphing both equations What is the solution y x 21 y 2x 30
What changes occur in the genes sex cell
The major source of primary socialization for children in the United States is
In a paragraph contrast the characteristics between anorexia and bulimia
Reducing individual purchases of disposable products would be a form of management of pollution prevention of pollution cure for pollution
What was The Grapes of Wrath based on A the Lost Generation authors B the Harlem Renaissance C the Dust Bowl D the Roaring 20s